# 3cm3cm2cm2cm

### v1

@@ 3 Quantification of significance of Odderon-exchange
The covariance matrix has been published together with the differential cross-section of elastic scattering at $\sqrt{s}=13$ TeV by TOTEM [57] and on this dataset, we have checked with a Levy series expansion method [19], that a fit with the full covariance matrix and another fit with our adopted PHENIX method gave the same minimum, the same central values for the fit parameters and the same errors of these parameters, within one standard deviation. This suggests that indeed the two methods are equivalent in our case too, and we utilized and adapted the PHENIX method for the comparision of two datasets, where the covariance matrix of at least one of these datasets was not publicly available.
At $\sqrt{s}=2.76$ TeV, in Ref. [15], the TOTEM Collaboration published the $pp$ differential cross section data with separated type A and type B errors. The source of the TOTEM $pp$ differential cross section data, measured at $\sqrt{s}=7$ TeV, is Ref. [51]. In addition, the values of $|t|$ were determined together with their errors of type A and B as given in Table 5 of Ref. [46] and Table 3 of Ref. [58]. The $t$-independent, type C errors cancel from the $H(x)$ scaling functions, as they multiply both the numerator and the denominator of $H(x)$. The sources of the TOTEM $pp$ differential cross section data, measured at $\sqrt{s}=8$ TeV, are Refs. [16, 11] with errors of type A and B. The second reference specifies data without horizontal errors, only the bin widths are given that significantly overestimate those errors. Therefore, we have performed calculations with more realistic but likely still well overestimated horizontal errors of half bin widths, as well, resulting in much higher significances. The D0 collaboration did not publish type B errors for its differential cross-section data at $\sqrt{s}=1.96$ TeV [35]. We have thus fixed the correlation coefficient of these D0 type B errors to zero. The input values of the nuclear slope parameters $B$ and the elastic cross sections $\sigma_{\rm el}$ are summarized in Table 1, together with the appropriate references.
@@ 5 Conclusion
Using the $\chi^{2}$ method, the combined significance of Odderon exchange grows from 6.26 to 7.08 $\sigma$, as shown in Table 3, as obtained from TOTEM data on elastic $pp$ scattering at 7 and 8 TeV and all D0 data on elasttic $p\bar{p}$ scattering at 1.96 TeV. As also indicated in Table 3, Stouffer’s method for the combination of significances yields similar results, resulting in a 7.10 $\sigma$ combined statistical significance of odderon exchange using all the D0 elastic $p\bar{p}$ data at $\sqrt{s}=1.96$ TeV, Ref. [37] and all the TOTEM data at $\sqrt{s}=7$ and $\sqrt{s}=8$ TeV, Refs. [46, 11]. Let us note, that the domain of validity of this $H(x,s|pp)$ scaling includes these three energies, but so far this property has been published only in a model-dependent way [44, 8]. The domain of validity of this $H(x|pp)$ scaling has been found to include $\sqrt{s}=1.96$, $2.76$, $7$ and $8$ TeV in a model-independent way as well as presented at various conferences [52], but these results go well beyond the scope of the present manuscript and will be written up in full details in a separate publication.
Acknowledgments. We acknowledge inspiring and useful discussions with W. Guryn, G. Gustafson, V.A. Khoze, E. Levin, L. Lönnblad, M. Strikman as well as with members of the D0 and TOTEM collaborations. R.P. is supported in part by the Swedish Research Council grant No. 2016-05996, by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 668679). T.Cs., T.N., A.S. and I.Sz. were partially supported by the NKIFH grants K-133046 and K-147557 as well as by the EFOP 3.6.1-16-2016-00001 grant (Hungary). I. Sz. was supported as well as by the ÚNKP-21-3 and ÚNKP-22-3 New National Excellence Program of the Hungarian Ministry for Innovation and Technology from the National Research, Development and Innovation Fund. Our collaboration has also been partially supported by the COST Action CA15213 (THOR) and by the MATE KKP 2023 &\&& 2024 grants.

### v3

@@ 3 Quantification of significance of Odderon-exchange
The left panel of Fig. 3 compares the $H(x)$ scaling function of elastic $pp$ collisions at $\sqrt{s}=8$ TeV to that of $p\bar{p}$ collisions at $\sqrt{s}=1.96$ TeV. In this case, adopting the method of Ref. [57], the confidence level of the agreement of the $H(x)$ scaling functions is found to be lower then $0.015$%, with a minimum of $\chi^{2}/{\rm NDF}=46.4/17$. Hence, the difference between these scaling functions is statistically significant, an at least 3.79 $\sigma$ effect, and it represents our result for the odderon observation from the comparison of the $\sqrt{s}=8$ TeV $pp$ and $\sqrt{s}=1.96$ TeV $p\bar{p}$ datasets in the $5\leq x\lessapprox 20$ acceptance domain. This seems to be a conservative and robust result as we find that this value can only increase for variations both in the procedure itself and in the $\chi^{2}$ definition.
@@
The covariance matrix has been published together with the differential cross-section of elastic scattering at $\sqrt{s}=13$ TeV by TOTEM [18], and on this dataset, we have checked with a Levy series expansion method [23], that a fit with the full covariance matrix and another fit with our adopted PHENIX method gave the same minimum, the same central values for the fit parameters and the same errors of these parameters, within one standard deviation. This suggests that indeed the two methods are equivalent in our case too, and we utilized and adapted the PHENIX method for the comparision of two datasets, where the covariance matrix of at least one of these datasets was not publicly available.
At $\sqrt{s}=2.76$ TeV, in Ref. [19], the TOTEM Collaboration published the $pp$ differential cross section data with separated type A and type B errors. The source of the TOTEM $pp$ differential cross section data, measured at $\sqrt{s}=7$ TeV, is Ref. [52]. In addition, the values of $|t|$ were determined together with their errors of type A and B as given in Table 5 of Ref. [52] and Table 3 of Ref. [52]. The $t$-independent, type C errors cancel from the $H(x)$ scaling functions, as they multiply both the numerator and the denominator of $H(x)$.
The D0 collaboration did not publish type B errors for its differential cross-section data at $\sqrt{s}=1.96$ TeV [14]. We have thus fixed the correlation coefficient of these D0 type B errors to zero. The input values of the nuclear slope parameters $B$ and the elastic cross sections $\sigma_{\rm el}$ are summarized in Table 1, together with the appropriate references. The sources of the TOTEM $pp$ differential cross section data, measured at $\sqrt{s}=8$ TeV, are Refs. [20, 15] with errors of type A and B. The second reference specifies data without horizontal errors, only the bin widths are given that significantly overestimate those errors. Therefore, we have performed calculations with more realistic but likely still well overestimated horizontal errors of half bin widths, as well, resulting in significances much higher as compared to the significances presented in Table  2.
@@ 5 Conclusion
With the new $\sqrt{s}$ = 8 TeV $pp$ data of the TOTEM collaboration we have further strengthened the observation of Odderon that we have already shown in our previous analysis with 7 TeV data. Using the $\chi^{2}$ method, the combined significance of Odderon exchange grows from 6.26 to 7.08 $\sigma$, as shown in Table 3, as obtained from TOTEM data on elastic $pp$ scattering at 7 and 8 TeV and all D0 data on elastic $p\bar{p}$ scattering at 1.96 TeV. As also indicated in Table 3, Stouffer’s method for the combination of significances yields similar results, resulting in a 7.10 $\sigma$ combined statistical significance of odderon exchange using all the D0 elastic $p\bar{p}$ data at $\sqrt{s}=1.96$ TeV, Ref. [14] and all the TOTEM data at $\sqrt{s}=7$ and $\sqrt{s}=8$ TeV, Refs. [52, 15]. Let us note, that the domain of validity of this $H(x,s|pp)$ scaling includes these three energies, but so far this property has been published only in a model-dependent way [46, 9]. The domain of validity of this $H(x|pp)$ scaling has been found to include $\sqrt{s}=1.96$, $2.76$, $7$ and $8$ TeV in a model-independent way as well as presented at various conferences [55], but these results go well beyond the scope of the present manuscript and will be written up in full details in a separate publication.
Acknowledgments. We acknowledge inspiring and useful discussions with W. Guryn, G. Gustafson, L. Jenkovszky, V.A. Khoze, E. Levin, L. Lönnblad, M. Strikman as well as with members of the D0 and TOTEM collaborations. R.P. is supported in part by the Swedish Research Council grant No. 2016-05996, by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 668679). T.Cs., T.N., A.S. and I.Sz. were partially supported by the NKIFH grants K-133046 and K-147557 as well as by the EFOP 3.6.1-16-2016-00001 grant (Hungary). I. Sz. was supported as well as by the ÚNKP-21-3 and ÚNKP-22-3 New National Excellence Program of the Hungarian Ministry for Innovation and Technology from the National Research, Development and Innovation Fund. Our collaboration has also been partially supported by the COST Action CA15213 (THOR) and by the MATE KKP 2023 &\&& 2024 grants.